CO2 (g) is placed in a sealed 5.0L vessel at 701 K and 10.0 atm. The vessel is heated to 1401K and the CO2 (g) partially decomposes to CO )g) and O2 (g). The final pressure in the vessel is 22.5 atm.

a. What is the complete balanced equation for this reaction?

b. How many moles of CO2 are present in the vessel initially (show work)

c. How many moles are present in the flask at the end of the reaction?

d. How many moles of CO2 CO and O2 are present in the flask after the reaction?

e. What is the partial pressures due to CO2, CO, and O2 in the flask at the end of the reaction?
Answer #1

Ideal gas eqt PV=nRT P- Pressure = 10.o atu V. Potal rolume 25.02 To Temperature. Folk R2 gan Constant - 8.314 Jk most. no no2 con 2 co & O2 dnitial 0.8698 Mole after rean 0. 8698-20 22 2. Total no. of mole = 0. 8698-24 + 2x tn = 0.8698 + x 0. 8698 +